A baseball player hits a ball so that it leaves the bat at a speed of 45 m/s at 60 degree above the horizontal and at a height of 2 meters above the ground. Find its position at t=3s.
Answer #1

horizontal position

Dx = (v cos x) t

Dx = 45 cos 60 * 3 = 67.5 m

Vertical position

Dy = ho + (v sin x) t - 0.5*gt^2

Dy = 2 + 45 sin 60 * 3 - 0.5* 9.8* 3^2

Dy = 74.81 m

========

Comment in case any doubt, will reply for sure.. Goodluck